sq.rt(3+4i)+squ.rt(3-4i)=????? please solve it as soon as possible.

Grade:12th pass

1 Answers

Arun
4 years ago
 
Dear student
 
sqrt (3+ 4i ) = 2 + i
 
sqrt (3 – 4i) = i – 2
 
hence on adding = 2i
 
Hope it helps
 
Regards
Arun (askIITians forum expert)